### v2

# Backward Joint Model for the Dynamic Prediction of Both Competing Risk and Longitudinal Outcomes

Abstract
Joint modeling is a useful approach to dynamic prediction of clinical outcomes using longitudinally measured predictors. When the outcomes are competing risk events, fitting the conventional shared random effects joint model often involves intensive computation, especially when multiple longitudinal biomarkers are be used as predictors, as is often desired in prediction problems. This paper proposes a new joint model for the dynamic prediction of competing risk outcomes. The model factorizes the likelihood into the distribution of the competing risks data and the distribution of longitudinal data given the competing risks data. It extends the basic idea of the recently published backward joint model (BJM) to the competing risk setting, and we call this model crBJM. This model also enables the prediction of future longitudinal data trajectories conditional on being at risk at a future time, a practically important problem that has not been studied in the statistical literature. The model fitting with the EM algorithm is efficient, stable and computationally fast, with a one-dimensional integral in the E-step and convex optimization for most parameters in the M-step, regardless of the number of longitudinal predictors. The model also comes with a consistent albeit less efficient estimation method that can be quickly implemented with standard software, ideal for model building and diagnostics. We study the numerical properties of the proposed method using simulations and illustrate its use in a chronic kidney disease study.
Wenhao Li1, Brad C. Astor2, Wei Yang3, Tom H. Greene4, Liang Li1,∗
1 Department of Biostatistics, The University of Texas MD Anderson Cancer Center, Houston, TX, U.S.A.
2 School of Medicine and Public Health, University of Wisconsin-Madison, Madison, WI, U.S.A.
3 Perelman School of Medicine, University of Pennsylvania, Philadelphia, PA, U.S.A.
4 School of Medicine, University of Utah, Salt Lake City, UT, U.S.A.
email: LLi15@mdanderson.org
Keywords: Competing risks; Dynamic prediction; Joint model; Multivariate longitudinal data; Survival analysis.
1 Introduction
Chronic kidney disease (CKD) is a major public health burden, affecting approximately 14% of the adult population in the US. It is characterized by a gradual and often irreversible deterioration of kidney function that may occur over years (Appel et al., 2023). The stages of CKD are defined based on kidney function, usually quantified by the glomerular filtration rate (GFR, in mL/min/1.73 m2). GFR estimates how well the kidneys filter waste and excess fluid from the blood. Stages 1, 2, 3A, 3B, 4 and 5 correspond to GFR in $\geqslant 90$, $60-89$, $45-59$, $30-44$, $15-29$, and $<15$ mL/min/1.73 m2, respectively. At the low end of stage 4 or stage 5, patients need to be prepared for entering End-Stage Renal Disease (ESRD), when either dialysis or kidney transplant is required to compensate for the malfunctioning kidney. More severe CKD is progressively associated with many adverse outcomes, including cardiovascular disease, stroke, heart failure, and peripheral vascular disease (Appel et al., 2023). Major causes of CKD in the US include diabetes mellitus and hypertension. CKD, aging, and these comorbidities may result in death, though the exact cause can be difficult to ascertain and is often multifactorial.
Predicting ESRD is important for CKD research, patient counseling, and clinical management of the disease. Various prediction models have been proposed (Lim et al., 2022), but they leave one or more of statistical challenges unaddressed. First, the prediction should ideally adapt to the ever-changing at-risk population and the potentially time-varying association between the predictor and outcome (Li et al., 2016). The at-risk population changes because of both aging and systematic differences between those who progress to ESRD early and those who do not. The relative importance of predictors may vary at various stages of the disease. Second, it is important to have the capacity to incorporate prior clinical history into prediction, as opposed to only the current health status. Third, the prediction of ESRD should adjust for the competing risk by death to avoid bias and produce proper interpretation. Fourth, not all patients are equally concerned about the near-future risk of ESRD. Those at earlier stages have low ESRD risk and are more interested in the future progression trajectory of GFR if ESRD or death does not occur (GFR is undefined when a patient reaches ESRD or death, obviously).
The statistical challenges above can be handled in the framework of dynamic prediction (Taylor et al., 2005; Yu et al., 2008; Rizopoulos, 2011). A typical approach is to estimate the conditional joint distribution of the longitudinal predictor variables (e.g., GFR) and survival data (e.g., the times of ESRD and death), given the time-invariant predictors. Once this joint distribution is estimated from the training dataset, the prediction for any subject in the validation dataset can be calculated straightforwardly from the conditional distribution given the longitudinal history before the prediction time. Therefore, the key to dynamic prediction is to build an appropriate joint model of longitudinal and survival data. Comprehensive reviews on this topic are available (Rizopoulos, 2012; Elashoff et al., 2016; Andrinopoulou et al., 2021), but the published statistical methods are insufficient to study the aforementioned CKD problem.
First, most published joint modeling research presented data applications and simulations with a very small number of longitudinal variables. The main reason is that the likelihood function of these models is complicated and contains an analytically intractable numerical integration for each of the $n$ subjects in the training dataset. Since the typical joint model uses shared random effects to quantify the association between longitudinal and survival data, the dimension of the integral increases linearly with the number of longitudinal variables, and the number of unknown parameters in the random effects distribution increases quadratically. In addition, the marginal log-likelihood function is usually not concave with multiple local maxima, which may cause non-convergence or computational instability in numerical optimization, especially when the number of unknown parameters is large. These are widely recognized difficulties in the implementation of such models. (Rizopoulos, 2012; Elashoff et al., 2016) They become particularly troublesome when we use joint models for prediction problems, which may require a large number of longitudinal predictors for improved predictive accuracy. Despite salient recent efforts to tackle this problem using numerical approximation to the high-dimensional integrals (Mauff et al., 2020; Rustand et al., 2022), it remains a challenge to design a computationally tractable, robust and unbiased method. Second, the aforementioned computational problem exacerbates the joint modeling with competing risk outcomes, whose multiple survival endpoints further increase the number of shared random effects. The limited literature on this topic presented numerical studies with a small number of random effects (or dimension of integrals), typically coming from one or two longitudinal variables, and with “brute force” maximum likelihood or Bayesian computation to tackle the model fitting problem (Elashoff et al., 2007a, b; Li et al., 2009, 2012; Andrinopoulou et al., 2014, 2015; Rajeswaran et al., 2018; Rizopoulos et al., 2023). When the number of longitudinal variables increases, it becomes increasingly difficult to use this strategy to maintain computational stability and avoid excessive computing time. Third, while there is previous literature on the prediction of future longitudinal biomarker trajectories (Taylor et al., 2005; Yu et al., 2008; Rizopoulos, 2012), this was done without conditioning on the fact that the future survival endpoint does not occur and hence is not suitable for our research problem (e.g., the predicted future GFR at year 3 is meaningless if the patient dies before that time).
In this paper, we propose a new joint model that avoids these limitations and is therefore suitable for the CKD research problem and beyond. It is built on the basic idea of the recently proposed backward joint model (BJM) (Shen and Li, 2021) in the sense that it uses a similar likelihood factorization of longitudinal and survival data and hence shares many desirable features, which we will review in Section 2. However, the model in this paper has three novel contributions. First, it accommodates competing risk outcomes. To distinguish it from the previous BJM (Shen and Li, 2021), we call this new model crBJM. Second, we use crBJM to predict future longitudinal biomarkers in the absence of future terminal events. This is a new statistical problem, which can be viewed as joint prediction of future longitudinal and future time-to-event data. Third, we study two formulations of crBJM that cover various censoring schemes. The crBJM model and estimation can be easily scaled for a large number of longitudinal variables because the dimension of the numerical integral is always capped at 1. For most model parameters, the log-likelihood has a concave shape and estimation has closed form solution, which facilitates convergence. The proposed method also comes with a simpler version that can be implemented quickly with standard statistical software to obtain consistent albeit slightly less efficient estimation. The proposed crBJM is also flexible in the sense that it can use various competing risk models or longitudinal data models as its building blocks for the prediction.
2 Data, Estimands and Models
2.1 Data and notation
The training dataset has $n$ subjects, indexed by $i=1,2,...,n$. Let $\tilde{T}_{i}$ denote the true time to a terminal event and $C_{i}$ denote the censoring time. All the time variables in this paper are expressed as the time gap since baseline unless stated otherwise. The terminal event can be one of $J$ event types, denoted by $\tilde{D}_{i}\in\{1,2,...,J\}$. In our motivating data application in CKD, $J=2$, and the terminal event is either ESRD ($j=1$) or death ($j=2$). The observed survival time is $T_{i}=\min(\tilde{T}_{i},C_{i})$. The observed event type is $D_{i}=\delta_{i}\tilde{D}_{i}\in\{0,1,...,J\}$, where the censoring indicator $\delta_{i}=\bm{1}(\tilde{T}_{i}<C_{i})$ equals $1$ when the terminal event is observed and 00 when censored. Let $\bm{V}_{i}$ denote the vector of baseline covariates. Suppose that there are $M$ longitudinally measured continuous biomarkers, denoted by $m=1,2,...,M$. We denote the repeated measures of the $m$-th biomarker of subject $i$ by $\bm{Y}_{mi}$, with the corresponding vector of time points $\bm{t}_{mi}$. The number of longitudinal biomarker measurements is $n_{mi}$, which varies by biomarker and subject. The measurement times of a biomarker may be irregularly spaced between baseline and $T_{i}$, and different biomarkers may be measured at different time points. Let $\bm{Y}_{i}=(\bm{Y}_{1i}^{T},\bm{Y}_{2i}^{T},...,\bm{Y}_{Mi}^{T})^{T}$ and $\bm{t}_{i}=(\bm{t}_{1i}^{T},\bm{t}_{2i}^{T},...,\bm{t}_{Mi}^{T})^{T}$ collect all the longitudinal biomarker data and their measurement times in vector forms. We follow two conventional assumptions from the joint modeling literature. First, the biomarker measurement times are non-informative in the sense that the distribution of $\bm{t}_{i}$ in $(0,T_{i})$ is independent of the rest of the data. Second, the censoring is independent in the sense that $C_{i}$ is conditionally independent of $\bm{Y}_{i}$ and $\tilde{T}_{i}$ given $\bm{V}_{i}$.
2.2 Prediction estimands
The models in this paper are motivated by the research objective, estimating the predicted distribution of the survival and longitudinal outcomes given the longitudinal history prior to the time of prediction. Therefore, we must define the estimands of this prediction problem before describing the models. We use the subscript $o$ to index a generic subject in the validation dataset. Let $s$ be the time of prediction and $\overline{\bm{Y}_{o}(s)}$ represent the longitudinal biomarker history prior to $s$ (i.e., all the biomarker measurements before $s$). $\Delta$ denotes a pre-specified prediction horizon. The prediction estimand of scientific interest is defined and derived in (1). This is the probability of having the terminal event of type $j$ ($j=1,2,...,J$) in the next $\Delta$ years (or other time units), given that the subject is still at-risk at the prediction time $s$ ($T_{o}>s$) and given the observed history ($\bm{V}_{o}$, $\overline{\bm{Y}_{o}(s)}$). The derivation used the independent censoring assumption.
$P(s<\tilde{T}_{o}\leqslant s+\Delta,\tilde{D}_{o}=j|\overline{\bm{Y}_{o}(s)},T_{o}>s,\bm{V}_{o})=\dfrac{P(\overline{\bm{Y}_{o}(s)},s<\tilde{T}_{o}\leqslant s+\Delta,\tilde{D}_{o}=j|\bm{V}_{o})}{P(\overline{\bm{Y}_{o}(s)},\tilde{T}_{o}>s|\bm{V}_{o})},~{}\forall j.$
We use $P(.)$ to denote probability or probability density, when there is no ambiguity. We are also interested in the dynamic prediction of future longitudinal biomarker distribution. This is formulated as the estimand in (2) for a future time $t>s$. The notation $\bm{Y}_{o}(t)$ denotes the generic subject $o$’s longitudinal biomarker value at $t$.
$P(\bm{Y}_{o}(t)=\bm{y}|\tilde{T}_{o}>t,\overline{\bm{Y}_{o}(s)},T_{o}>s,\bm{V}_{o})=\frac{P(\bm{Y}_{o}(t)=\bm{y},\overline{\bm{Y}_{o}(s)},\tilde{T}_{o}>t|\bm{V}_{o})}{P(\overline{\bm{Y}_{o}(s)},\tilde{T}_{o}>t|\bm{V}_{o})},$ (2)
This is the probability density function of the longitudinal biomarker at a future time $t$, given that the subject is at-risk by the time of prediction ($T_{o}>s$), and the observed history $\{\overline{\bm{Y}_{o}(s)},\bm{V}_{o}\}$, and conditional on the absence of the terminal event by time $t$. A key novel aspect of this definition is that the predicted longitudinal biomarker at a future time $t$ can only be defined when the terminal event does not occur by that time. In our data application, for example, the GFR does not exist if the patient has reached ESRD or death. We may choose $t=s+\Delta$ to obtain the biomarker prediction at the horizon, or choose $t$ to be any time between $s$ and $s+\Delta$ to obtain the predicted longitudinal trajectory in $(s,s+\Delta)$. The notation in (2) is designed for the most general case that we predict the multivariate density of all the $M$ biomarkers. In our data application, we are interested in the single most important longitudinal biomarker and similar expression can be formulated for that situation. Given the estimated density function, various interpretable features of the density, such as quantiles, mean and mode, can be subsequently obtained, as illustrated in the data application.
2.3 Backward joint models with extrapolation or two parts
If the conditional joint distribution $P(\bm{Y}_{o},\tilde{T}_{o},\tilde{D}_{o}|\bm{V}_{o})$ can be estimated from the training dataset, then the prediction estimands in (1) and (2) can be easily calculated. Since $P(\bm{Y}_{o},\tilde{T}_{o},\tilde{D}_{o}|\bm{V}_{o})=P(\tilde{T}_{o},\tilde{D}_{o}|\bm{V}_{o})P(\bm{Y}_{o}|\tilde{T}_{o},\tilde{D}_{o},\bm{V}_{o})$, we propose sub-models for the two conditionals. The first is a competing risk regression with baseline covariates. The second sub-model can be a multivariate linear mixed model when $\bm{Y}_{o}$ is continuous. Since $\bm{Y}_{o}$ occurs prior to $\tilde{T}_{o}$, we call this approach a competing risk “backward” joint model (crBJM). This likelihood decomposition resembles the pattern mixture model in the missing data literature but there are fundamental differences in model formulation, estimation and scope of application, as elucidated previously in backward joint model without competing risk (BJM) (Shen and Li, 2021). From (1) and (2), all that is needed for prediction purposes is the conditional joint distribution $P(\bm{Y}_{o},\tilde{T}_{o},\tilde{D}_{o}|\bm{V}_{o})$. In fact, any model that leads to the estimation of this joint distribution can be used here. The model does not have to reflect the data-generating mechanism in the sense that it consists of the conditional distributions of later data given earlier data. The data-generating mechanism of real data is usually far more complicated than a statistical model can describe.
[…]
A sub-model for survival data can identify both $P(\tilde{T}_{o}=u|\bm{V}_{o})$, $t<u\leqslant\tau_{max}$, and $P(\tilde{T}_{o}>\tau_{max}|\bm{V}_{o})$. Two sub-models for longitudinal data are needed to identify $P(\overline{\bm{Y}_{o}(s)}|\tilde{T}_{o}=u,\bm{V}_{o})$, $t<u\leqslant\tau_{max}$ and $P(\overline{\bm{Y}_{o}(s)}|\tilde{T}_{o}>\tau_{max},\bm{V}_{o})$ for those with $\tilde{T}>\tau_{max}$. These subjects with $\tilde{T}>\tau_{max}$ are referred to as long-term survivors (LTS; Wang et al, 2020(Wang et al., 2020)). In real data analysis, we can empirically use the estimated marginal distribution of $\tilde{T}$ to decide whether the extra longitudinal sub-model for the LTS is needed. If the Kaplan-Meier curve of $\tilde{T}$ descends to near 00, that suggests that the LTS group is likely very small. We call it the crBJM with extrapolation (crBJM-EX). The extrapolation means that the estimated distribution $P(\tilde{T}|\bm{V})$ extrapolates beyond the support of the data sample. Alternatively, the LTS sub-model is needed and we call it the two-part crBJM (crBJM-TP). For crBJM-TP, the prediction time $s<\tau_{max}-\Delta$ because there is no observed terminal event beyond $\tau_{max}$ and the model does not extrapolate beyond the data range. For crBJM-EX, there is no such restriction because the model extrapolates.
2.4 Sub-models for competing risk outcomes
Any competing risk regression model with time-invariant covariates can be fitted to $\{T_{i},D_{i},\bm{V}_{i};i=1,2,...,n\}$ and estimate $P(\tilde{T}_{o},\tilde{D}_{o}|\bm{V}_{o})$. Common choices include the cause-specific hazard models, the Fine-Gray sub-distribution hazard model, or the vertical modeling (Nicolaie et al., 2010), among others. We used vertical modeling in the numerical studies because it directly models $P(\tilde{T}_{i}|\bm{V}_{i})$, which is convenient for predicting longitudinal data in (2). We compared the crBJMs with vertical modeling and with cause-specific hazard models, and the predictive performances were similar in our data application.
The likelihood of a vertical model for $P(\tilde{T}_{i},\tilde{D}_{i}|\bm{V}_{i})$ can be factorized as the likelihood product of $P(\tilde{T}_{i}|\bm{V}_{i})$ and $P(\tilde{D}_{i}|\tilde{T}_{i},\bm{V}_{i})$. The former can be formulated as a semi-parametric Cox model, a parametric Weibull model, or any other survival regression model that fits the data well. The latter is a multinomial regression model, according to Nicolaie et al.(Nicolaie et al., 2010). Specifically, $P(\tilde{D}_{i}=j|\tilde{T}_{i}=t,\bm{V}_{i})=\dfrac{\exp(\bm{B}(t)^{T}\bm{\xi}^{(\tilde{T})}_{j}+\bm{V}_{i}^{T}\bm{\xi}^{(\bm{V})}_{j})}{\sum_{j=1}^{J}\exp(\bm{B}(t)^{T}\bm{\xi}^{(\tilde{T})}_{j}+\bm{V}_{i}^{T}\bm{\xi}^{(\bm{V})}_{j})},j=1,...,J.$. The $\bm{B}(t)$ is a vector of pre-specified regression spline basis to accommodate the nonlinear effect of $\tilde{T}_{i}$ on the event types. When there are two event types, this is a logistic regression with partially linear specification for the covariates $\tilde{T}_{i}$ and $\bm{V}_{i}$. The estimation of vertical modeling follows Nicolaie et al(Nicolaie et al., 2010).
2.5 Sub-models for longitudinal outcomes
For the crBJM-EX, $P(\bm{Y}_{o}|\tilde{T}_{o},\tilde{D}_{o},\bm{V}_{o})$ is modeled by a multivariate linear mixed model, with $T_{i}$, $D_{i}$, $\bm{V}_{i}$, $\bm{t}_{i}$ and their interactions as fixed-effect covariates and with random effects (intercept, or intercept and slope) to account for the longitudinal correlation and the correlation among different biomarkers. The sub-model of the $m$-th longitudinal biomarker ($m=1,2,...,M$) can be formulated in the following general form.
$Y_{mi}(t)=\sum_{l=0}^{L}\vartheta_{ml}(\bm{V}_{i},\phi(\tilde{T}_{i}),\tilde{D}_{i})g_{l}(t)+\bm{u}_{mi}(t)^{T}\bm{b}_{mi}+\epsilon_{mi}(t)~{}~{},~{}~{}t\in[0,\tilde{T}_{i}).$ (4)
$Y_{mi}(t)$ is subject $i$’s $m$-th biomarker, measured at time $t$. In the training data, $t\in\bm{t}_{mi}$. The first term describes the conditional mean trajectory (CMT), given baseline covariates ($\bm{V}$), terminal event time ($\tilde{T}$) and type ($\tilde{D}$). The CMT is defined for subjects with the same trajectory lengths ($\tilde{T}$), because it does not lead to proper interpretation to average the longitudinal biomarker trajectories of different lengths when the follow-up duration is correlated with longitudinal data, a defining feature of joint modeling problems that is sometimes called informative dropout. The second term includes random effects that describe the subject-specific, mean-zero deviation from the CMT. The random intercept ($\bm{u}_{mi}(t)=1$), random intercept plus slope ($\bm{u}_{mi}(t)=(1,t)^{T}$), or other nonlinear random effect terms can be used. The $g_{l}(t),l=0,1,...,L$ are basis functions that characterize the shape of the CMT. For example, the linear CMT has $L=1$, $g_{0}(t)=1$, and $g_{1}(t)=t$. The coefficients of the basis functions, $\vartheta_{ml}(\bm{V}_{i},g_{1}(\tilde{T}_{i}),\tilde{D}_{i})$, are interpreted as the intercept, slope, etc., allowing the CMT to vary with $\bm{V}_{i}$, $\tilde{T}_{i}$ and $\tilde{D}_{i}$. For example, $\vartheta_{m0}(\bm{V}_{i},\phi(\tilde{T}_{i}),\tilde{D}_{i})=\bm{\beta}_{m1}^{T}\bm{V}_{i}+\nu_{m0}\tilde{D}_{i}+\nu_{m2}\phi(\tilde{T}_{i})+\nu_{m4}\phi(\tilde{T}_{i})\tilde{D}_{i}$, $\vartheta_{m1}(\bm{V}_{i},\tilde{T}_{i},\tilde{D}_{i})=\bm{\beta}_{m2}^{T}\bm{V}_{i}+\nu_{m1}\tilde{D}_{i}+\nu_{m3}\phi(\tilde{T}_{i})+\nu_{m5}\phi(\tilde{T}_{i})\tilde{D}_{i}$. The pre-specified transformation $\phi(.)$, e.g. $\log$, may be used to reduce the skewness of the time variable $\tilde{T}_{i}$.
The residuals $\varepsilon_{mi}(t)$ are independent noises with mean zero and variance $\sigma_{e,m}^{2}$. Extensions to residuals with auto-correlation are straightforward. The random effects from all $M$ biomarkers $\bm{b}_{i}=(\bm{b}_{1i}^{T},\bm{b}_{2i}^{T},...,\bm{b}_{Mi}^{T})^{T}$ follows a multivariate normal distribution with mean zero and covariance matrix $\bm{\Omega}$. When model (4) has random intercept and slope, $\bm{\Omega}$ is a $2M\times 2M$ matrix that is not restricted to be block diagonal, allowing the biomarkers to be correlated.
The crBJM-TP has an extra longitudinal sub-model for the LTS:
$Y_{mi}(t)=\sum_{l=0}^{L}\vartheta_{ml}^{e}(\bm{V}_{i})g_{l}(t)+\bm{u}_{mi}^{e}(t)^{T}\bm{b}^{e}_{mi}+\epsilon_{mi}^{e}(t)~{}~{},~{}~{}t\in[0,\tau_{max})~{}\mathrm{and}~{}\tilde{T}_{i}>\tau_{max}.$ (5)
The superscript $e$ symbolizes quantities in the “extra” LTS model. The model formulation and interpretation are similar to (4) but with a few differences. First, model (5) does not involve $\tilde{T}_{i}$ and $\tilde{D}_{i}$ in the covariates because the LTS includes all subjects who survive beyond $\tau_{max}$ regardless of their actual $\tilde{T}_{i}$ and $\tilde{D}_{i}$. Second, since the LTS is a heterogeneous group, models (5) and (4) cannot have the same random effect variance or residual variance. Presumably, the variances are larger in model (5).
3 Estimation
The crBJM’s likelihood factorization $P(\bm{Y}_{o},\tilde{T}_{o},\tilde{D}_{o}|\bm{V}_{o})=P(\tilde{T}_{o},\tilde{D}_{o}|\bm{V}_{o})P(\bm{Y}_{o}|\tilde{T}_{o},\tilde{D}_{o},\bm{V}_{o})$ suggests a two-step estimation procedure. In the first step, we fit the competing risk regression model (Section 2.4). Many standard software packages are available. In the second step, we estimate $P(\bm{Y}_{o}|\tilde{T}_{o},\tilde{D}_{o},\bm{V}_{o})$ from the longitudinal sub-models in Section 2.5 and this is the focus of this section. We propose two approaches. Section 3.1 presents a simple method that can be conveniently implemented with standard linear mixed model software. This method only uses training data subjects whose terminal event times are observed and those still at-risk by $\tau_{max}$. Since many censored data are not used, we call this approach the complete-case analysis (CCA). A remarkably useful feature of the crBJM is that the CCA produces consistent estimation, albeit less efficient. In contrast, typical joint models of longitudinal and survival data would require specialized software and they have limited functionality and flexibility. The CCA of crBJM enables quick and convenient exploratory analysis, model building, model assumption diagnosis, and preliminary prediction accuracy assessment. Once the model specification is completed, we can use the fully likelihood-based analysis in Section 3.2 to obtain efficient estimators and the final results of the prediction model and prediction accuracy.
3.1 Complete-case analysis (CCA)
To see why the parameters of longitudinal sub-model can be consistently estimated with the complete cases in the training data (i.e., $D_{i}\neq 0$), we observe that, for any $j=1,2,...,J$, $P(\bm{Y}|\tilde{T}=t,\tilde{D}=j,\bm{V})=P(\bm{Y}|\tilde{T}=t,C>t,\tilde{D}=j,\bm{V})=P(\bm{Y}|T=t,D=j,\bm{V})$. The simple derivation is justified by the conditional independence of censoring. It holds at any $t\in(0,\infty)$ for the crBJM-EX, and any $t\in(0,\tau_{max})$ for the crBJM-TP. $P(\bm{Y}|T=t,D=j,\bm{V})$ can be estimated via a linear mixed model according to the specification in (4). Additionally, $P(\bm{Y}|\tilde{T}>\tau_{max},\bm{V})=P(\bm{Y}|\tilde{T}>\tau_{max},C>\tau_{max},\bm{V})=P(\bm{Y}|T>\tau_{max},\bm{V})$ in the crBJM-TP. Hence, model (5) can be estimated from the subset of training data with $T_{i}>\tau_{max}$. Consistent estimation of CCA has been demonstrated in previous research on models with a similar likelihood factorization (Li et al., 2018; Wang et al., 2020; Shen and Li, 2021; Wang et al., 2023).
3.2 EM algorithm for the analysis of all data
This section describes the EM algorithm that calculates the maximum likelihood estimators (MLE) of the parameters in the longitudinal sub-models. Since it uses both censored and non-censored data, the estimation is more efficient than the CCA (Li et al., 2018; Wang et al., 2020; Shen and Li, 2021). The censored $\tilde{T}_{i}$ and $\tilde{D}_{i}$ are treated as missing data in the complete data log-likelihood. Since the parameters in the competing risk sub-model are estimated consistently in Section 2.4 and are fixed at their estimated values during the EM iterations, the method in this section is a type of pseudo maximum likelihood estimation (Gong and Samaniego, 1981).
Let $\bm{\Theta}$ denote the parameters of the longitudinal sub-model. With crBJM-TP, $\bm{\Theta}$ includes additionally the parameters $\bm{\Theta}^{e}$ from the LTS sub-model (5). The estimated parameters from the survival sub-model are denoted by $\hat{\bm{\Theta}}_{surv}$. Since $\hat{\bm{\Theta}}_{surv}$ is fixed throughout the EM iterations, we omit it for the formula below. In the following, we mainly describe the EM algorithm of the crBJM-EX, but will point out the changes needed when the crBJM-TP is used. The complete data log-likelihood is $\sum_{i=1}^{n}\left\{\log P(\bm{Y}_{i}|\tilde{T}_{i},\tilde{D}_{i},\bm{V}_{i};\bm{\Theta})+\log P(\tilde{T}_{i},\tilde{D}_{i}|\bm{V}_{i})\right\}$. The observed data of all subjects in the training dataset is collected in $\{\bm{Y},\bm{T},\bm{\delta},\bm{D},\bm{V}\}$. The EM algorithm uses the consistent estimators from the CCA as the initial values in the first EM iteration. At the beginning of the $q$-th iteration ($q=1,2,...$), the estimated parameter is denoted by $\hat{\bm{\Theta}}^{(q)}$.
In the E-step of the $q$-th iteration, we calculate the conditional expectation of the complete data log-likelihood as in (6). This expectation is taken over the conditional distribution of the missing data (censored $\tilde{T}_{i}$ and $\tilde{D}_{i}$) given the observed data, with parameters in this distribution set at $\hat{\bm{\Theta}}^{(q)}$.
$~{}\sum_{i=1}^{n}E_{(\tilde{T}_{i},\tilde{D}_{i})}\left\{\log P(\bm{Y}_{i}|\tilde{T}_{i},\tilde{D}_{i},\bm{V}_{i};\bm{\Theta}^{(q+1)})+\log P(\tilde{T}_{i},\tilde{D}_{i}|\bm{V}_{i})\Big{|}\bm{Y},\bm{T},\bm{\delta},\bm{D};\hat{\bm{\Theta}}^{(q)}\right\}$ (6)
[…]
There is always a one-dimensional integral in (6), regardless of how many longitudinal biomarker variables are involved. This integral is needed only for censored subjects. In contrast, the widely used shared random effects joint model (Rizopoulos, 2012) has $n$ integrals in its likelihood, and the dimension of the integral increases with the number of longitudinal variables. This feature is the key reason that crBJM is computationally much easier. The integration in (6) goes from $T_{i}$ to $\infty$. With crBJM-TP, this calculation is split into two parts, analogously to (3). The first integrates from $T_{i}$ to $\tau_{max}$ using model (4) and the second uses the LTS model (5).
The integral in (6) can be calculated with the trapezoidal rule. If $T_{i}$ is a censoring time, we assume that $\tilde{T}_{i}$ falls in one of a series of pre-specified connected intervals, denoted by $\{\mathbb{L}_{ik}=[a_{ik},b_{ik});k=1,...,K_{i}\}$. The intervals are connected (e.g., $b_{ik}=a_{i(k+1)}$) and starts from $T_{i}$ ($a_{i1}=T_{i}$). With crBJM-EX, we can set the ending time of the last interval at a very large survival time (e.g., 100 year old). With crBJM-TP, the last interval is $[\tau_{max},\infty)$ and the LTS model is used on this interval. A moderate $K_{i}$ can produce accurate results when the residual survival beyond $T_{i}$ is not very large or when the integrand is not excessively wiggly (Li et al., 2018). The target function (6) of crBJM-EX then becomes:
$\sum_{i=1}^{n}\sum_{k=1}^{K_{i}}\sum_{j=1}^{J}\left\{\log P(\bm{Y}_{i}|\tilde{T}_{i}=\dfrac{a_{ik}+b_{ik}}{2},\tilde{D}_{i}=j,\bm{V}_{i};\bm{\Theta}^{(q+1)})\right\}P(\tilde{T}_{i}\in\mathbb{L}_{ik},\tilde{D}_{i}=j|\bm{Y},\bm{T},\bm{\delta},\bm{D};\hat{\bm{\Theta}}^{(q)}).$ (7)
With crBJM-TP, the target function is the same except that the terms corresponding to the last interval ($\mathbb{L}_{iK_{i}}$) are replaced by $\left\{\log P(\bm{Y}_{i}|\tilde{T}_{i}\in\mathbb{L}_{iK_{i}},\bm{V}_{i};\bm{\Theta}^{(q+1)})\right\}P(\tilde{T}_{i}\in\mathbb{L}_{iK_{i}}|\bm{Y},\bm{T},\bm{\delta},\bm{D};\hat{\bm{\Theta}}^{(q)})$.
In the M-step, (7) is maximized with respect to $\bm{\Theta}^{(q+1)}$, producing an updated estimator $\hat{\bm{\Theta}}^{(q+1)}$ that will enter the next ($q+1$-th) iteration. The only terms containing the unknown parameters are $\log P(\bm{Y}_{i}|\tilde{T}_{i}=(a_{ik}+b_{ik})/2,\tilde{D}_{i}=j,\bm{V}_{i};\bm{\Theta}^{(q+1)})$. These are log of multivariate normal density functions, which are quadratic forms of the regression coefficients in the longitudinal sub-models. $P(\tilde{T}_{i}\in\mathbb{L}_{ik},\tilde{D}_{i}=j|\bm{Y},\bm{T},\bm{\delta},\bm{D};\hat{\bm{\Theta}}^{(q)})$ is calculated in the E-step and treated as fixed in the maximization of the M-step. Therefore, the regression coefficients have closed-form solutions (the variance-covariance matrices of the random effects do not; see below). This feature, in conjunction with using the consistent estimators from the CCA as the initial value of the EM algorithm, speeds up the computation considerably and improves the robustness of the iterations. We have observed a 100% convergence rate with the EM and the computing time is much shorter compared to the shared random effects joint model.
The number of regression coefficients in the crBJM increases linearly with the number of longitudinal biomarkers. However, the number of parameters in the variance-covariance matrix of the random effects goes up quadratically, which poses a challenge for any joint models that aim to accommodate multivariate longitudinal data. However, the crBJM has a special feature that helps avoid this computational burden. During the EM algorithm, we fix the variance-covariance matrix of the random effects and residuals at their consistently estimated values from the CCA. This approach is justified by the pseudo maximum likelihood estimation (Gong and Samaniego, 1981).
3.3 Variance estimation
We use the bootstrap for variance estimation. Specifically, we randomly sample the subjects in the training data with replacement, creating a bootstrap dataset. The bootstrap dataset is put through the crBJM estimation procedure to produce point estimators of the model parameters. This process is repeated 200 times, and the variances of the estimated model parameters can be calculated as the sample variance of the 200 point estimators. For a subject in the validation dataset, the prediction (1) or (2) is a function of the estimated crBJM model parameters and the subject’s data prior to the prediction time. The variance of the prediction can be calculated by Delta-method to account for the sampling variability of the estimated crBJM model. As an alternative approach that avoids the mathematical derivation of the Delta-method, we can also calculate the prediction using each of the 200 estimated model parameters from the bootstrap. The resulting sample variance quantifies the variance of the prediction. Since the estimated model parameters are approximately normal, a Wald-type confidence interval can be obtained from the estimated variance.
3.4 Prediction accuracy measures
We follow the previous literature in choosing the prediction accuracy measures. For predicting competing risk outcomes, we use the time-dependent Brier score (BS) and the time-dependent AUC (Definition A in Wu and Li, 2018(Wu and Li, 2018)), both designed for competing risk data. For predicting the future longitudinal biomarkers, we use RMSE (root mean squared error). Since our data application is CKD research, we also use two commonly used measures in that field for predicting GFR, the P30 and P50 (Levey, 1999). They are the proportions of subjects for whom the predicted GFR falls within 30% or 50% of the observed GFR at the horizon. Higher AUC, lower BS, lower RMSE, higher P30 and P50 suggest better predictive accuracy. When the GFR measurements are irregularly spaced and there is no GFR data at the horizon, the closest measurement prior to the horizon is used.
4 Data Application
We illustrate the proposed crBJM methodology by analyzing data from the African American Study of Kidney Disease and Hypertension (AASK) Cohort Study. AASK is a longitudinal cohort study of 1,094 CKD subjects. Detailed descriptions of AASK data were reported previously (Li et al., 2016; Wu et al., 2020). By the end of the follow-up, 318 subjects (29%) reached ESRD and 176 subjects (16%) died before reaching ESRD. We did not consider survival after ESRD in a semi-competing risk analysis because the patient experience and treatments were substantially different after reaching ESRD. We used two baseline predictors, age and gender. We considered three longitudinally measured predictive biomarkers, the estimated glomerular filtration rate (eGFR) using the MDRD equation, 24-hour urine protein to creatinine ratio (UPCr), and serum albumin. These biomarkers have known clinical relationships with the survival outcomes and hence are ideal choices as predictors.
Figure 2 displays the Kaplan-Meier curve of the composite endpoint of ESRD or death, whichever occurs first, and the cumulative incidence functions of ESRD and death separately. The distribution of the composite event in this population can be estimated up to 12 years, with more than 40% of the events expected beyond that time. This supported the use of crBJM-TP. For comparison purposes, we also presented results from crBJM-EX. We set $\tau_{max}=12$ years, the maximum observed event time in the data. We experimented with both vertical modeling and cause-specific hazard modeling as the survival sub-model, and they produced similar prediction accuracy. The reported results here came from crBJM with vertical modeling. Figure 3 displays the CMTs of eGFR, which are approximately linear and suggest a clear association with the time of the composite endpoint and event types, with an interaction between the two.
Including more predictive longitudinal biomarkers improves the prediction accuracy. The crBJM remains computationally convenient in this situation. To illustrate the benefit of using more longitudinal predictors, we considered crBJMs with one (crBJM-1, with eGFR), two (crBJM-2, with eGFR and UPCr), or three (crBJM-3, with eGFR, UPCr, and albumin) longitudinal biomarkers. Our EM algorithm achieved 100% convergence rates in these analyses. Since this paper is mainly about prediction, we focus on prediction results here, and defer the extensive details on model specification, parameter estimation and interpretation, and sensitivity analysis to the Web Appendix. With crBJM-3, the extrapolation model (EX) has 48 parameters and the two-part model (TP) has 72 parameters. The crBJM includes a competing risk regression model and a linear mixed model for each longitudinal parameter. These are conventional statistical models in which all the parameters are well identified. The number of parameters characterizing the correlation of multiple longitudinal biomarkers increases quadratically with the number of biomarkers, but these parameters are inevitable in any joint models with multivariate longitudinal biomarkers. The TP model has 50% more parameters because it uses two parts to improve flexibility without extrapolation, but all the parameters in its LTS sub-model are identified.
We used five-fold cross-validation to evaluate the prediction accuracy at a horizon of 3 years (Table 1). Additional results under 1-year horizon are reported in the Web Appendix. In predicting ESRD and death, Table 1 shows that crBJMs with more longitudinal biomarkers generally performed better than those with less, with crBJM-3-TP performing similarly to or better than all other methods in both AUC and BS. The TP methods generally performed better than the EX methods, likely because it is less prone to model misspecification when the tail distribution of the composite event cannot be estimated. The performance of crBJM-3-EX rose on par with their counterpart TP model (crBJM-3-TP), while crBJM-2-EX and crBJM-1-EX did not. The performance of EX models is more difficult to speculate about probably because the accuracy of the extrapolation is unknown and cannot be verified from data. Nonetheless, crBJM-3-EX achieved similar predictive accuracy for ESRD and death as crBJM-3-TP did but with fewer parameters, making it an empirically useful model for prediction purposes. In predicting the eGFR, TP methods substantially outperformed EX methods. The BS appears to be more sensitive to between-method differences than the AUC, possibly because the latter is rank-based. The incidence of ESRD can be predicted with very high AUC because the treatment decision of dialysis or renal transplantation is strongly influenced by the patient’s eGFR.
For comparison purposes, we fitted the widely-used shared random effects joint model (SJM) using the JMBayes2 package in R. The model included all three longitudinal biomarkers (i.e., denoted by SJM-3 in Table 1), whose trajectories were modeled by linear mixed models with subject-specific random intercepts and random slopes, and these random effects were shared with the cause-specific hazard models as covariates. The crBJMs outperformed the SJM in predicting ESRD and death. The results of eGFR prediction are mixed. However, it seems that the SJM does not condition on the absence of the composite event when predicting eGFR (Section 7.2 of Rizopoulos, 2012), which makes this comparison difficult. The crBJM is computationally much faster than the SJM. To complete the five-fold cross-validation on a personal computer with 2.9 GHz CPU and 32 GB RAM, SJM-3 took 3 hours whereas crBJM-3-TP with EM algorithm took 13 minutes. The CCA of crBJM can implemented with standard statistical software and the computing time is almost negligible. The computing time will multiply if bootstrap is used to derive the variance of the estimated parameters. However, variance estimation is not as essential in prediction problems as in treatment effect estimation. The guaranteed convergence of the crBJM estimation ensures that the bootstrap iterations are robust.
Figure 4 depicts individual dynamic predictions of ESRD and death for two AASK patients. The patient on the left (A) had stable eGFR before year 3 but went into a steady decline from year 3 to year 8. The UPCr also increased over time, indicating increasing kidney damage. The predicted future probability of ESRD increases sharply in response. This patient eventually reached ESRD. The patient on the right had stable biomarkers and the predicted ESRD risk is low. The mortality risk, however, increases over time, presumably due to both aging and low ESRD risk (a competing risk phenomenon). Figure 5 illustrates two useful graphical presentations of the joint prediction of competing risk and longitudinal outcomes. The first uses the quantile curves to visualize how the predicted eGFR distribution conditional on absence of ESRD or death varies over time, in addition to the predicted cumulative incidence functions of ESRD and death. This patient’s eGFR was within a stable range between the two prediction times, and the predicted ESRD was similar as a result. However, we can observe a subtle decrease in the predicted eGFR in the second prediction, perhaps due to the lower observed eGFR just prior to the prediction and increased UPCr. The cumulative incidence of death increases due to time elapsing and aging. The second graphical presentation is clinically more interpretable (Hu et al., 2012), where the future probabilities of being in each CKD stage and death are visualized. From year 2 to year 4, the eGFR stabilized around 30, which triggered the model to lower the probability of ESRD but enlarge the probability area corresponding to CKD stage 4 (eGFR between 15 and 29). The probability of death increased slightly due to time elapsing and aging.
We performed a sensitivity analysis to study whether the prediction accuracy is affected by the specifications for random effects in (4), random intercept only or both random intercept and slope, and the choices of $\phi(.)$ in that model, $\phi(t)=\log(t)$ or $\phi(t)=t$. The random slope adds flexibility to the individual-level longitudinal trajectories within each survival stratum (i.e., given $\tilde{T}$) but also adds more parameters to the variance matrix for the random effects. A log transformation may fit the data better when $\tilde{T}$ is skewed. Web Table S4 shows that the prediction accuracy is not sensitive to these alternative formulations. We speculate that the survival time in the AASK data is not very skewed. For this reason, we used the simpler model specification of random intercept and no log transformation.
5 Simulation
We conducted simulations to study the finite sample performance of the proposed estimation procedure for both TP and EX models, and to compare the relative efficiency of CCA vs. EM algorithm. The CCA and EM methods both produce consistent estimation, but CCA is easier to use, making it worthwhile to assess its efficiency cost. We already compared the predictive performance of crBJM and SJM in the AASK data. We do not compare them in the simulations because they operate under different modeling assumptions and it is impossible to simulate data so that both are correctly specified. Within the likelihood framework, consistent and efficient estimation of the crBJM guarantees optimal predictive accuracy.
We conducted and reported simulation studies for the crBJM-EX and crBJM-TP separately. In each study, we simulated 300 independent Monte Carlo datasets of $n$ subjects from the crBJM-EX model (or crBJM-TP model), $n=300$ or $1,000$, analyzed each dataset using both the CCA and EM algorithm, and calculated the bias of estimated model parameters and the relative efficiency between CCA and EM. We simulated one baseline covariate and three longitudinal biomarkers. The longitudinal sub-models had a linear CMT and random intercept. The competing risk model is a vertical model with a Weibull regression for the time to composite event. The censoring rate is 40% in the crBJM-EX data. For crBJM-TP, the overall censoring rate is 40% and 20% subjects were censored at $\tau_{max}$. The crBJM-EX and crBJM-TP were fit to their respective datasets so that we can evaluate the estimation methods under the correct model specification. The CCA was used to obtain the initial values for the EM. All EM algorithms converged under the criterion that all longitudinal sub-model parameters differed by no more than $10^{-4}$ in consecutive iterations. Further details of the data-generating schemes are in the Web Appendix. Figure 1 shows that the finite-sample biases of the EM estimators are generally small and drop to nearly zero when the sample size increases from 300 to 1,000, supporting consistent estimation. The EM algorithm has notable efficiency improvement over the CCA, especially among the LTS sub-model parameters, probably because the CCA only uses the subjects at-risk by $\tau_{max}$ to estimate the LTS, which could have a small sample size.
6 Discussion
The idea of factorizing the joint likelihood into survival distribution and conditional distribution of longitudinal data and survival data is not new and has been used in various statistical problems. See the literature reviews in Li et al (2018)(Li et al., 2018) and Shen & Li (2021). (Shen and Li, 2021) Shen & Li (2021)(Shen and Li, 2021) appears to be the first to use this idea in dynamic prediction problems. This paper is novel in several aspects: (1) predicting competing risk outcomes, (2) dual approaches (EX and TP) designed for different scenarios, and (3) predicting the future distribution of longitudinal outcome conditional on being at risk. Our proposed estimand (2) is novel and should be used when the survival outcome is a terminal event in the sense that the longitudinal outcome is undefined after its occurrence, such as the ESRD or death in our data application. Since future longitudinal and future survival data are correlated and our prediction algorithm fully accounts for that correlation, this paper can be viewed as the first in the literature to tackle this dynamic “joint” prediction problem, with a joint model of both longitudinal and survival (competing risk) data. Of note, another popular method for dynamic prediction, the landmark modeling, would find difficulty for dynamic joint prediction problems, as discussed below.
The estimation algorithm of crBJM ensures that model fitting remains computationally tractable with a large number of longitudinal variables. This is highly desirable in prediction problems, but the widely used SJM can be computationally challenging in such situations. We compared the prediction accuracy of crBJM and SJM in our data application, and crBJM performed notably better. While a generalizable conclusion cannot usually be drawn from one data analysis, and it is difficult to compare the prediction accuracy in simulations because crBJM and SJM operate under different modeling assumptions, two features of the crBJM help improve its predictive performance. The first is model formulation. The SJM was proposed decades ago for explaining the association between longitudinal and survival data, or for informative dropout problems. It was recently repurposed for use in dynamic prediction problems (Rizopoulos, 2011). In contrast, the crBJM was designed from the ground up for dynamic prediction. As shown in Section 2, we first proposed the predictive estimand in (1) and (2) and then proposed a sub-model for each component in that estimand. Any appropriate sub-models for those components can be used in combination to improve prediction accuracy in various contexts, an interesting topic for future research. The second is model estimation. The SJM must be estimated by specialized software, which often offers limited flexibility. The lengthy computing time may also limit the scope of exploration. The crBJM is much faster to estimate, and the model formulation can even be explored in CCA with standard software. We believe that the ease of estimation accelerates the development of crBJM variants for various research problems.
Another approach for dynamic prediction is the landmark modeling (van Houwelingen and Putter, 2011; Zheng and Heagerty, 2005). The joint modeling has many advantages over landmark modeling, including efficient likelihood-based estimation and easily incorporating the longitudinal history into the prediction. However, the landmark modeling is easy to implement, often with standard software, and it can handle large number of predictor variables (Ferrer et al., 2019; Li et al., 2023). With crBJM, we have made joint modeling as computationally efficient, scalable, and easy to develop as landmark modeling. In addition, we are not aware of any landmark modeling approach that can perform a dynamic joint prediction as defined above. This is not straightforward or convenient within the landmark modeling framework because it would require modeling the joint distribution of residual lifetime (possibly with competing risks) and future longitudinal data at various landmark times, and aggregating models over different landmark times.
A limitation of this paper is that the longitudinal sub-models are parametric. The crBJM models the CMT, not the individual-level trajectories. The CMT usually has less curvature and hence a parametric model may suffice. Model checking can also be done conveniently using CCA. Nonetheless, it is difficult to ensure that many longitudinal biomarkers are all modeled adequately with parametric models. Future work is needed to improve the flexibility of the fixed effect part with semi-parametric or non-parametric specification. The random effect part can also be improved by using penalized splines for subject-specific trajectory deviation from the CMT, though that would cause a substantial increase in the number of parameters in the covariance matrix of spline coefficients. However, this difficulty challenges all joint models with multivariate nonlinear longitudinal trajectories. Recent work suggested that, even if the random effect part is misspecified, the rest of the crBJM model can be estimated consistently (Wang et al., 2023). In the sensitivity analysis (Web Table S4), we found that adding a random slope causes little change in the prediction accuracy. Since the crBJM characterizes the association between survival outcome and longitudinal biomarkers through the fixed effects (as opposed to using random effects, such as in the SJM), it remains to be studied whether the crBJM is robust to misspecification of the random effect part.
Another limitation of the crBJM is that it sometimes requires an extra LTS sub-model. SJM does not have this need because it uses a time-varying hazard function. While the LTS can be easily estimated, it increases the number of parameters by about 50%. At a chosen $\tau_{max}$, the percent of subjects contributing to the LTS estimation is $P(\tilde{T}>\tau_{max})P(C>\tau_{max})$. The censoring distribution may depend on staggered study entry, administrative censoring, and dropout patterns. When $P(C>\tau_{max})$ is low, we may need to decrease $\tau_{max}$ to ensure enough sample size for the LTS estimation. The longitudinal and survival data beyond $\tau_{max}$ are artificially censored as a result. Another limitation is that this paper assumes that the biomarker measurement times are non-informative in the sense that they are not related to the longitudinal or competing risk data. This assumption may not hold in clinical data. We are unaware of any dynamic prediction methods designed to work with informative observational times. This issue warrants further research.
The crBJM and the previous BJM(Shen and Li, 2021) use the independent censoring assumption that the censoring time is conditionally dependent of both terminal event and longitudinal data given the baseline covariates. This is also a standard assumption in the SJM literature. (Wulfsohn and Tsiatis, 1997; Diggle et al., 2009; Tsiatis and Davidian, 2004) However, the SJM can be extended to allow the censoring time to be correlated with both the terminal event and longitudinal data by adding a model for the censoring time given the shared random effects (Diggle et al., 2009), though this is rarely done in the SJM literature. Among longitudinal cohort studies collecting both longitudinal and survival data, independent censoring is perhaps more common than dependent censoring because the standard survival analysis relating survival outcomes to baseline variables is often used as the primary analysis, and joint modeling of longitudinal and survival data is among the secondary analyses (Tsiatis and Davidian, 2004). The independent censoring assumption is usually indispensable in standard survival analysis. Nonetheless, it is of future research interest to augment the backward joint models with a censoring time model and accommodate dependence between the censoring time and the longitudinal and terminal events data for some data applications.
Acknowledgements
This research is supported in part by NIH grants P30CA016672 and R01DK118079.
Supporting Information
Web Appendices, Tables, and Figures are available with this paper. Supplement A includes three sections. In Supplement A.1, we report additional details of the crBJM and its estimation. In Supplement A.2, we present additional details on the simulation study design. In Supplement A.3, we provide additional data analysis results from AASK.